OpenStudy (anonymous):

need help with this hard problem... (4,-6) write equation in point-slope form.....

OpenStudy (anonymous):

@CashBoy27 or @Jamierox4ev3r

OpenStudy (anonymous):

Rise over run

OpenStudy (anonymous):

huh???

OpenStudy (anonymous):

Point slope form is y=mx+b. You have x=4 and y=-6.

OpenStudy (anonymous):

would X=4 and Y= 6 my answer....

OpenStudy (anonymous):

No you need to put those answers into point slope form.

OpenStudy (anonymous):

ohhh ok..this is confusing... :(

OpenStudy (anonymous):

What do u mean.

OpenStudy (anonymous):

Take your time.

OpenStudy (anonymous):

y-4 = \[\frac{ 3 }{ 5 }\] (x + 6)

OpenStudy (anonymous):

Is this multiple choice?

OpenStudy (anonymous):

no sorry I meant that as the whole like answer..

OpenStudy (anonymous):

I got that. Where did you get 3/5?

OpenStudy (anonymous):

that's part of the equation sir

OpenStudy (anonymous):

Please don't call me sir. You posted that x=4 and y=-6. So you put it into the form y(which is -6)=mx(x is 4)+b. You don't have a b.

OpenStudy (anonymous):

sorry sir!!!!!

OpenStudy (anonymous):

A. \[y-4 = \frac{ 3 }{ 5 } \left( x + 6 \right) \] B. \[y-6 \frac{ 3 }{ 5 } \left( x + 4 \right) \\] C. \[y + 6 \frac{ 3 }{ 5 } \left( x-4 \right)\] D. \[y + 6 = \frac{ 3 }{ 5 } \left( x-4 \right)\]

OpenStudy (anonymous):

@Jamierox4ev3r need your help please